OpenStudy (anonymous):

PLEASE I NEED HELP!! A right triangle has base (x - 5) units and height (4x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points)

OpenStudy (anonymous):

I only need part c, i already got part a and b...

Parth (parthkohli):

\[\text{Area of triangle}= {\rm \frac{1}{2}\times base \times height = ~}\frac{1}{2} \times (x - 5) \times (4x-10) = \cdots\]

Parth (parthkohli):

The closure property of polynomials says that the product of two polynomials is also a polynomial.

OpenStudy (anonymous):

so that would be my answer for part c?? @ParthKohli

Parth (parthkohli):

No, but that certainly is a good explanation of what's happening here.

OpenStudy (anonymous):

ok @ParthKohli so thats part of my answer for part c because its asking to explain the closure property of multiplication of polynomials and you just explained it to me, but I also have to explain in part c using the solution obtained in part b.

OpenStudy (anonymous):

anyway I figured it out already thanks for the help @ParthKohli

Parth (parthkohli):

Hmm, you're welcome.